the length of a rectangle is 4 less than twice its width. if the area of the rectangle is 20, find the width of the rectangle to the nearest tenth
Answer by checkley75(3666)

L=2W-4 IS THE LENGTH WHILE W=WIDTH
AREA=W*L
20=W(2W-4)
20=2W^2-4W
2W^2-4W-20=0
USING THE QUADRATIC EQUATION WE GET
W=(4+-SQRT[-4^2-4*2*-20])/2*2
W=(4+-SQRT[16+160])/4
W=(4+-SQRT176)/4
W=(4+-13.2665)/4
W=(4+13.2665)/4
W=17.2665/4
W=4.3166 ANSWER FOR THE WIDTH.
L=2*4.3166-4
L=8.6312-4
L=4.6312 ANSWER FOR THE LENGTH.
PROOF
4.3166*4.6312=20
20=20